Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

pm1weekly_2004-115_2004-121

Mission Operations Days: 2004/115 to 2004/121
April 24, 2004 00:00:00 GMT to April 30, 2004 00:00:00 GMT


Aqua Spacecraft and MODIS Instrument Status:

PM-1 "Aqua" is in Fine Pointing Normal Mode
FM1 MODIS is in B-Side Nominal Science Mode
ComponentState(s)Comment
BlackbodyA Off; B On (285K)Nominal Science Mode
Calibration ElectronicsA Off; B OnNominal Science Mode
Control ProcessorA Off; B OnNominal Science Mode
Nadir DoorUnlatched, OpenNominal Science Mode
Space View DoorUnlatched, OpenNominal Science Mode
Solar Diffuser DoorUnlatched, ClosedNominal Science Mode
FDDI FormatterA Off; B OnNominal Science Mode
FDDI PortA Off; B OnNominal Science Mode
FIFO Memory#1 & #2 Off; #3 & #4 OnNominal Science Mode
Format ProcessorA Off; B OnNominal Science Mode
Power Supplies#1 Off, #2 OnNominal Science Mode
PV VIS FPAA Off; B OnNominal Science Mode
PV NIR FPAA Off; B OnNominal Science Mode
PV SMIR FPAA Off; B OnNominal Science Mode
PV LWIR FPAA Off; B OnNominal Science Mode
PC LWIR FPAA Off; B OnNominal Science Mode
RC Outgas HeatersAll 3 (CS, IS, OS) OffNominal Science Mode
RC LWIR FPA HeaterOffNominal Science Mode
RC SMIR FPA HeaterOn (83K)Nominal Science Mode
Scan AssemblyA Off; B OnNominal Science Mode
SDSMOffNominal Science Mode
SRCAOffNominal Science Mode
PS1 Survival HeatersEnabledNominal Science Mode
PS2 Survival HeatersEnabledNominal Science Mode
Timing GeneratorA Off; B OnNominal Science Mode
Flight SoftwareRev BJ + 1 set of patches
  • CP07 to Safe Mode - No FR OFF Command (1 patch)
Nominal Science Mode
Inhibit Ids SetNoneNominal Science Mode

This Week's Completed MODIS Activities:

Saturday, April 24, 2004

2004/115 00:00:00 MCL MOD_BB_TEMP_SET_285K_B.1 BB Temp set to 285K for CP B Coefficients
2004/115 03:00:00 MCL MOD_BB_TEMP_SET_290K_B.1 BB Temp set to 290K for CP B Coefficients
2004/115 07:00:00 MCL MOD_BB_TEMP_SET_315K_B.1 BB Temp set to 315K for CP B Coefficients
2004/115 07:00:30 MCL MOD_BB_DCYCLE_FULL.1 Set blackbody heater duty cycle to FULL
2004/115 11:00:00 MCL MOD_BB_TEMP_SET_270K_B.1 BB Temp set to 270K for CP B Coefficients
2004/115 11:00:30 MCL MOD_BB_DCYCLE_THIRD.1 Set blackbody heater duty cycle to THIRD

Sunday, April 25, 2004

2004/116 07:00:00 MCL MOD_BB_TEMP_SET_290K_B.1 BB Temp set to 290K for CP B Coefficients

Monday, April 26, 2004

2004/117 21:19:03-21:33:01 MCL MOD_OA16_SDSM_SCREEN.2 OA-16: SD/SDSM Screened - B Side
2004/117 22:57:56-23:11:54 MCL MOD_OA15_SDSM_OPEN.1 OA-15: SD/SDSM Open - B Side

Tuesday, April 27, 2004

2004/118 01:28:37-01:31:08 MCL MOD_OA27_PV_ECAL.1 OA-27: Photovoltaic bands electronics calibration
2004/118 04:40:00 MCL MOD_SRCA_SETUP_NO_L2.2 SRCA setup activity using lamp #4 in lieu of lamp #2
2004/118 04:43:29-05:35:11 MCL MOD_OA22_SRCA_FULL_SPECT_1.1 OA-22: SRCA Full Spectral Calibration Part 1 30W-lamp data collection is from 04:59:23-05:35:10
2004/118 06:39:14-07:13:46 MCL MOD_OA22_SRCA_FULL_SPECT_2.1 OA-22: SRCA Full Spectral Calibration Part 2 30W-lamp data collection is from 06:39:14-07:13:45
2004/118 08:00:00 MCL MOD_SRCA_SETUP_NO_L2.2 SRCA setup activity using lamp #4 in lieu of lamp #2
2004/118 08:01:17-08:52:18 MCL MOD_OA22_SRCA_FULL_SPECT_3.1 OA-22: SRCA Full Spectral Calibration Part 3 10W-lamp data collection is from 08:17:44-08:52:17
2004/118 09:54:49-10:32:50 MCL MOD_OA22_SRCA_FULL_SPECT_4.1 OA-22: SRCA Full Spectral Calibration Part 4 10W-lamp data collection is from 09:54:49-10:32:49

Wednesday, April 28, 2004

2004/119 23:30:05-23:30:23 Real-Time MOD_SET_BB_TEMP(1733) BB Temp set to 285K for CP B Coefficients

Thursday, April 29, 2004

None

Friday, April 30, 2004

Lunar Roll Maneuver (-6.9795 degree roll)
2004/121 16:54:31 MCL MOD_OA03_FR_NIGHT_RATE (03:23 early)
2004/121 17:04:14 Real-Time Aqua FOT commands S/C to low-fidelity pointing
2004/121 17:12:30-17:19:16 MCL MOD_OA08_ROLL_MAN
2004/121 17:30:32 Real-Time Aqua FOT commands S/C to high-fidelity pointing
2004/121 17:50:43 MCL MOD_OA03_FR_DAY_RATE (03:23 late)

This Week's Scheduled MODIS Activities Not Completed:

None

Upcoming MODIS Events:

Calibrations for May will include biweekly SDSM, monthly SRCA Full Radiometric, and bimonthly SRCA Full Spatial calibrations.

Maneuvers:

The next Aqua MODIS Lunar Calibration Roll will be scheduled for 2004/151 (May 30, 2004). Data collect time is 04:27:14 to 04:33:44 GMT. The roll angle will be -7.8218 degrees.

Aqua performed a Drag Make-Up (DMU) Maneuver on 2004/112 (April 21, 2004). The burn was 34 seconds long and performed at 15:49:00 GMT.

No official date has yet been scheduled for the Aqua Deep Space Calibration pitch maneuver(s).

MODIS Anomalies:

None

General Instrument Comments:

FM1 MODIS is in Science Mode. The SVD and NAD are Open, the survival heaters are enabled, and the Black Body is heated to 285 K.

MODIS Telemetry Trends:

The FM1 MODIS instrument telemetry is trending nominally.

Non-MODIS Significant Events:

The EMOS mission status webpage (which includes instrument data loss records) can be found at the following URL. Note that a password is now required to access the EOS mission links webpages. There are directions at the URL below for how to obtain a username/password.

http://jupiter02.gsfc.nasa.gov/links/links.htm

Limited Life Item Status:

SRCA 10W Lamp #1: 232.2 of 500 hours
SRCA 10W Lamp #2: 184.3 of 500 hours
SRCA 10W Lamp #3: 197.0 of 500 hours
SRCA 10W Lamp #4: 70.2 of 500 hours

SRCA 1W Lamp #1: 507.3 of 5000 hours
SRCA 1W Lamp #2: 269.8 of 5000 hours

Solar Diffuser Door: 2200 of 3022 Movements
Nadir Aperture Door: 1053 of 1316 Movements
Space View Door: 632 of 1316 Movements